Three years ago, Mary was three times as old as Ann. Three years hence she will be twice as old Find the present age of eac.
Answer by ankor@dixie-net.com(22740)

let m = Mary's present age
let a = Ann's
:
Write an equation for each statement, simplify
:
Three years ago, Mary was three times as old as Ann.
m - 3 = 3(a - 3)
m - 3 = 3a - 9
m = 3a - 9 + 3
m = 3a - 6
Three years hence she will be twice as old
m + 3 = 2(a + 3)
m + 3 = 2a + 6
m = 2a + 6 - 3
m = 2a + 3
:
m = m, therefore we can write the equation
3a - 6 = 2a + 3
3a - 2a = 3 + 6
a = 9 yrs old is Ann
:
Find m using m = 2a + 3
m = 2(9) + 3
m = 21 yrs is Mary
:
:
Confirm our solution in the first statement equation
21 - 3 = 3(9 - 3)
18 = 3(6)